OpenStudy (anonymous):

simplify each expression. show work. 2 + 6 x 8 divided by 4

OpenStudy (igreen):

\(\dfrac{2 + 6 \times 8}{4}\) \(\dfrac{2 + 48}{4}\) \(\dfrac{50}{4}\) \(12.5\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

i acgree with @iGreen please make sure to reward him/her